OpenStudy (anonymous):

An alloy of tin is 15% tin and weighs 20 pounds. A second alloy is 10% tin. How many pounds of the second alloy must be added to the first to get a 12% mixture?

OpenStudy (nurali):

Let x pound of the 2nd alloy is added to make the final composition 12%. (20 X 0.15 + 0.1 x ) / (20 + x ) = 12 / 100. Solving this gives x = 30. That is 30 pound of the second alloy when added to 20 lb of the first alloy, after homogenisation will give an alloy of 12% final composition. Total weight being 50 lb.